State the value of the discriminant. Then determine the number of real roots of the equation.
n(8n + 10) = -15

Answer by stanbon(75887)

State the value of the discriminant. Then determine the number of real roots of the equation.
-----
discriminant = b^2 - 4ac
--------
n(8n + 10) = -15
8n^2 + 10n + 15 = 0
----
a = 8 ; b = 10 ; c = 15
b^2-4ac = 10^2 - 4*8*15 = 100 - 480 = -380
---------------
Ans: Since the discriminant is negative, the number of Real roots is zero.
Cheers,
Stan H.
-------------